OpenStudy (anonymous):

What value of p makes the equation true? -9.3 = 3p A) -27.9 B) -12.3 C) -6.3 D) -3.1

OpenStudy (alexandervonhumboldt2):

divide both sides by 3: 3p/3=-9.3/3 p=-9.3/3 p=-3.1

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Don't forget to close your question! ☺